OpenStudy (anonymous):

i have a question.. so My teacher gave me a problem like this 15x+5=0 her answer: 5(x+3)=0 x+3=0

OpenStudy (anonymous):

are you solving for x?

OpenStudy (anonymous):

sorry i was not done but the answet is X=-3

OpenStudy (anonymous):

yeah, what is your question? :)

OpenStudy (anonymous):

-7x-14=0

OpenStudy (anonymous):

-7x-14=0 +14 -7x=-14 ___ ___ -7 -7 x=2 :)

OpenStudy (anonymous):

then a nother problem is Factor: 15x+5 she got as a asnwer 5(3x+1 my question is how do u solve this one 18+24X

OpenStudy (anonymous):

3(6+8x)

OpenStudy (anonymous):

HOW did u get that? O.o

OpenStudy (anonymous):

find the greatest common factor (GCF)... which is 3 for this problem

OpenStudy (anonymous):

wait can u explain more im sorry ...

OpenStudy (anonymous):

then you divide.. 18/3=6 24/3=8 = 3(6+8x)

OpenStudy (anonymous):

use distributive property to check, and you get the original answer.

OpenStudy (anonymous):

ok thank you (:

OpenStudy (anonymous):

how do you solve this 5x^2+15x=0